OpenStudy (anonymous):

MEDAL SOLVE (cos Θ − cos Θ)2 + (cos Θ + cos Θ)2

zepdrix (zepdrix):

Hey there :) Unless your teacher wants you to do this a specific way, notice that you have a cosine minus cosine, which gives you zero there. And in the other brackets, simply combine your like-terms.\[\Large\rm (\cos \theta-\cos \theta)^2+(\cos \theta+\cos \theta)^2\]\[\Large\rm (0)^2+(2\cos \theta)^2\]

zepdrix (zepdrix):

And then square both the 2, and the cosine.

OpenStudy (anonymous):

0 +2cos

zepdrix (zepdrix):

No you still need to square the 2cos(theta).\[\Large\rm (2\cos \theta)^2=2^2\cos^2\theta\]

OpenStudy (anonymous):

8

OpenStudy (anonymous):

it is 4cos^2 theta

zepdrix (zepdrix):

yes.